# Med-R3: Enhancing Medical Retrieval-Augmented Reasoning of LLMs via Progressive Reinforcement Learning

### v3

@@ B.2 Reward Modeling Details
We provide an in-depth description of the reward computation process associated with progressive RL described in the main content. As shown in the right part of Figure˜6, during the verification phase of each training stage, we first parse the model’s natural language output into a structured knowledge graph representation in the form of Equation˜1. The stage-specific reward metric is then applied to assess both the natural language response and the structural and logical accuracy of the reasoning trajectory encoded in the knowledge graph.
B.2.1 Stage 1: Reasoner Cultivation
@@
The primary objective of this stage is twofold: (1) to regularize the model’s output format, ensuring structured and consistent responses, and (2) to cultivate its ability to perform logical reasoning when answering medical questions. In addition, by incorporating reward signals based on the number of knowledge retrievals, we encourage the model to make increased use of external medical knowledge sources, which lays the groundwork for enhancing the model’s retrieval capabilities in the subsequent stage. The reward is composed of four aspects: format, answer correctness, reasoning process and retrieval number, as depicted in Figure˜6. While format and answer correctness are commonly used reward components, we place particular emphasis on the rationale behind the design of rewards for reasoning process and retrieval number.
@@
  • Alignment with High-Quality Reasoning Trajectories: We aim to guide the model’s reasoning process toward high-quality trajectories observed in frontier models. To this end, our evaluation criteria incorporate semantic and structural characteristics of reference reasoning paths (see Equation˜4 to Equation˜7), enabling the model to learn from expert-like inference patterns.
  • Capturing Medical Reasoning Specifics: Medical reasoning often requires both logical coherence and comprehensive coverage of relevant entities and their relationships. Unlike mathematical or programming tasks that emphasize rigid logical order, medical reasoning allows for multiple valid pathways, provided they are semantically rich and clinically grounded Wu et al. (2023); Gao et al. (2025). Accordingly, we introduce: (1) a structural coherence score based on the sequential logic of the reasoning path, as outlined in Equation˜7, and (2) a completeness score evaluating the coverage of key entities and relations, as shown in Equation˜6.
  • Robust Evaluation via Maximum Similarity Matching: For each training instance, we provide multiple reference reasoning trajectories. However, these references may not be highly similar to one another due to the diversity of valid clinical reasoning. Therefore, in designing Equation˜5 to Equation˜7, we opt to use the maximum similarity score between the model’s reasoning process and the reference reasoning processes, both in terms of semantics and structure, rather than an average score, to ensure robust and meaningful scoring.
  • Flexibility and Generalization: Leveraging the strengths of reinforcement learning (RL), we avoid rigidly constraining the model’s reasoning trajectory. Instead, we encourage it to approximate high-quality reasoning behaviors while allowing room for variation and creativity. This flexible guidance enhances the model’s generalization capability when faced with novel or out-of-distribution reasoning tasks. To support this goal, we design a multi-dimensional reward scheme that captures various aspects of medical logic reasoning, as depicted in Equation˜4.
As detailed in Algorithm˜1, we have also provided the algorithm implementation of $R_{logical}$ defined in Equation˜7.
@@
Table 5: Comparison of average retrieval frequency with and without $\mathcal{R}_{retrieval\_num}$ in Stage 1.
Retrieval Number.  One of the key motivations for incorporating this reward in Stage 1 is to lay the foundation for enhancing the model’s retrieval capabilities in the subsequent stage. Specifically, only when the model develops the habit of querying external knowledge upon encountering unfamiliar concepts during reasoning, can there be meaningful room for optimizing its retrieval behavior. Our experiments also confirm this intuition: without such a reward signal in Stage 1, certain models tend to rely solely on internal knowledge and cease performing external retrievals altogether. As observed in Table˜5, using Qwen2.5-7B as an example, the average number of knowledge retrievals during logical reasoning drops significantly when the retrieval count component is omitted from the reward function.
B.2.2 Stage 2: Retriever Awakening
@@
The core reward metric in this phase is the retrieval effectiveness, which evaluates both the quality of the retrieved documents and their influence within the model’s overall reasoning trajectory, as outlined in Equation˜9. For document quality assessment, we adopt the Evidence-Based Medicine (EBM) grading framework illustrated in Figure˜5 to classify the retrieved documents into different quality levels. To evaluate the contribution of these retrieved documents to the model’s reasoning trajectory, we compute the ratio of retrieved knowledge graph triples relative to the total number of triples in the reasoning trajectory, as defined in Equation˜11. This metric reflects the extent to which external knowledge contributes to the logical inference process.
B.2.3 Stage 3: Dual-Process Collaboration
After the previous two stages which have separately enhanced the model’s reasoning and retrieval cabilities, in the final training stage, the model’s end-to-end (E2E) performance on medical problem-solving is directly optimized based on the accuracy of the final answers. As a result, the reward function in this stage consists solely of format and answer correctness.
@@ Appendix C Experiment Details
C.1 Baselines
In this section, we provide a comprehensive overview of the various models as well as methods that serve as baselines in our comparative analysis.
  • Close-Sourced Models: Close-sourced models are regarded as embodying the current peak performance across various capabilities of LLMs, and serve as the strongest baselines. Here we have selected GPT-4o-mini Hurst et al. (2024) for comparison.
  • Open-Sourced Medical-Specific Models: These models refer to domain-specific models that were trained specifically on medical data. We have selected MEDITRON-7B, MEDITRON-70B Chen et al. (2023), UltraMedical3-8B, and UltraMedical3.1-8B Zhang et al. (2024) to represent the open-sourced medical-specific models for comparison to assess Med-R3’s relative advantage. Specifically, the backbone model of UltraMedical3.1-8B is LLaMA3.1-8B-Instruct, which facilitates a more intuitive comparison.
  • Open-Sourced Medical Reasoning Models: Inspired by the breakthrough of OpenAI o1, these models are capable of complex medical reasoning, applying o1-like methods to the medical field. We have selected HuatuoGPT-o1-8B Chen et al. (2024b), MedS3-8B Jiang et al. (2025) and AlphaMed-8B Liu et al. (2025b) as competitors. HuatuoGPT-o1-8B improved models capabilities in medical reasoning by performing a two-stage SFT + RL training on verifiable medical problems. MedS3-8B equips the model with a self-evolution paradigm and proposes PRM-guided Vote-Sum (P-VS) strategy during inference to enhance long-chain reasoning capabilities in the medical domain. AlphaMed-8B is trained exclusively on rule-based RL. The backbone models of these competitors are all LLaMA-3.1-8B-Instruct, allowing for a more direct comparison with LLaMA-3.1-8B-Instruct + Med-R3.
@@ C.1 Baselines
  • Supervised Fine-Tuning (SFT): The model undergoes SFT using our constructed training dataset, with reasoning trajectories produced by frontier models (e.g., DeepSeek-V3), enriched with knowledge obtained through retrieval mechanisms. During training, retrieved documents are with masked losses.
  • General Retrieval-Augmented Reasoning RL: We also compare our method with approaches that enhance the model’s retrieval-augmented reasoning capabilities through RL in the general domain. R1-Searcher Song et al. (2025) is a two-stage outcome-based RL method designed to enhance the model’s capabilities of searching and integrating additional knowledge during the reasoning process. To ensure a more equitable comparison, we replaced the training dataset and the knowledge corpus for retrieval with one that matches our experimental setup. In the first stage, we trained the model for 1 epoch, and in the second stage for 2 epochs, with a total number of training samples matching that of Med-R3. ReSearch Chen et al. (2025) considers search operations as integral components of the reasoning chain, and trains models to reason with search via RL without using any supervised data on reasoning steps. Here we also train the model for 3 epochs to align with our experimental configuration. The RL process is implemented by the Group Relative Policy Optimization (GRPO) Shao et al. (2024) algorithm.
We evaluate these baselines under different inference methods, including Naive (direct generation), CoT (pure reasoning), RAG (pure retrieval), and CoT-RAG (interleaved reasoning and retrieval), as detailed in Table˜6.
Model
Training
Method
Inference
Method
MedQA-USMLE MedQA-MCMLE MedMCQA RareArena-RDC RareArena-RDS MMLU-Med NEJMQA MedXpertQA Avg.
@@
LLaMA3.1-8B
-Instruct
Naive 31.16 41.45 30.02 41.85 22.16 37.12 50.41 14.90 33.63
@@
Table 6: Comprehensive comparison of Med-R3 with baselines, where we have provided the inference strategy for each competitor. Here we have selected the highest-performing one (marked with ) to represent the optimal performance of each method, as summarized in Table˜1. * denotes our re-implementation with the same amount of our constructed training data for a fair comparison. represents the PRM guided Vote Sum (P-VS) strategy during inference, which unlocks the full potential of MedS3-8B Jiang et al. (2025). The best and second best of each model are in bold and underlined.
C.2 Benchmarks
To evaluate the performance of our proposed Med-R3 in both standard and real-world clinical scenarios of the medical domain, we have selected seven medical datasets, including the MedQA-USMLE, MedQA-MCMLE Jin et al. (2020), MedMCQA Pal et al. (2022), MMLU-Med Hendrycks et al. (2021), RareArena THUMedInfo (2025), MedXpertQA Alonso et al. (2024) and NEJMQA Katz et al. (2024). Among these benchmarks, MedQA series, MedMCQA and MMLU-Med are standard question-answering tasks, while RareArena, MedXpertQA and NEJMQA are related to real-world clinical scenarios. We employ LLM-as-Judge based on the frontier model DeepSeek-V3 Liu et al. (2024) to assess the correctness of the responses, then compute the accuracy scores as our evaluation metric.
We have introduced within-domain tasks in the section of training datasets details, where we use the training subsets for the RL process and test subsets (if no ground-truth labels provided, we use the development subsets) for in-domain evaluation. In the following, we give descriptions of the remaining out-of domain benchmarks:
  • MMLU-Med Hendrycks et al. (2021): The MMLU-Med task is composed of six medical domains, anatomy, clinical knowledge, professional medicine, human genetics, college medicine, and college biology, which is extracted from the MMLU benchmark.
  • NEJMQA Katz et al. (2024): The NEJMQA dataset is constructed from clinical case challenges sourced from The Lancet and the New England Journal of Medicine, which is centered on diagnostic reasoning using patient symptom information.
  • MedXpertQA Alonso et al. (2024): The MedXpertQA dataset incorporates specialty-specific evaluations and realistic clinical case questions derived from authentic medical practice. Here we utilize the text part of this benchmark.
C.3 More Experimental Observations
The performance of each baseline reported in Table˜1 in the main content represents the best result achieved in four different inference strategies. The comprehensive scores across all inference strategies are provided in Table˜6. In addition, we use the radar chart to visualize the performance of the methods that share the same backbone model of LLaMA3.1-8B-Instruct, as shown in Figure˜7. Furthermore, we summarize addition observations and insights derived from our experiments as follows:
@@ C.3 More Experimental Observations
RL boosts generalizable medical performances. We compare the performance of models after SFT and RL phase. Specifically, R1-Searcher, ReSearch, and Med-R3 all employ RL-based training approaches. As shown in Table˜6, on the in-domain benchmarks, the performance gap between SFT and R1-Searcher/ReSearch is relatively small, with SFT underperforming by 0.59% and 5.39%, respectively. However, on out-of-domain tasks, particularly on MMLU-Med, the gap becomes much larger, with SFT lagging behind R1-Searcher and ReSearch by 9.21% and 10.07%, respectively. A primary reason for this discrepancy lies in the nature of SFT as a strongly supervised learning paradigm. Under this setting, the model tends to learn by directly imitating the given reasoning trajectories, effectively memorizing problem-solving patterns rather than truly learning the underlying reasoning process. Consequently, its generalization capability is notably limited when faced with novel or unseen scenarios. It is worth noting that in our SFT setup, the training data includes retrieval-augmented reasoning trajectories that incorporate externally retrieved knowledge, which mitigates the inherent limitations of SFT to some extent. Nevertheless, a clear performance gap still remains.
Figure 7: Comparison of Med-R3 with baselines that utilize the LLaMA3.1-8B-Instruct as the backbone model.
C.4 Ablation Study and Analysis Details
We report the original performance scores of the models on each benchmark during training stage ablation as well as reward design ablation, as depicted in Table˜7 and Table˜8.
@@ C.4 Ablation Study and Analysis Details
1 $\rightarrow$ 2 $\rightarrow$ 3
Qwen2.5-7B 68.64 67.53 68.97 63.02 45.76 75.81 58.54 14.98 57.91
@@ 1 $\rightarrow$ 2 $\rightarrow$ 3
Table 7: Original scores for each benchmark of the ablation study on multiple training stages and sequential order. Order: sequential order of Stage 1, 2, and 3. Specifically, “1 &\& 2 &\& 3” refers to a joint training configuration in which reward functions from all three stages are merged and optimized concurrently. The best and second best scores of each model are in bold and underlined. We evaluate these models under the inference strategy of CoT-RAG (interleaved reasoning and retrieval).
@@
Table 8: Original scores for each benchmark of the ablation study on reward design. w/o: removing the reward component from the original metric. The best and second best scores are marked in bold and underlined. We employ Qwen2.5-7B for experimental analysis, and evaluate the model under the inference strategy of CoT-RAG (interleaved reasoning and retrieval).
C.4.1 The Role of Each Stage
Here we provide the detailed observations of Section˜4.2 in the main content.
Removing Stage 1: Stage 1 aims to enhance the reasoning capabilities of the models when solving medical problems. As depicted in Table˜2 (2 $\rightarrow$ 3), the absence of Stage 1 leads to a notable degradation of 7.30% in overall model performance. This reduction occurs because Stage 1 serves as the foundation for Stage 2. When the model lacks strong medical reasoning capabilities, it is unlikely to recognize when external knowledge is needed or to construct effective retrieval queries for obtaining supportive information. Consequently, the potential benefits offered by subsequent training stages are substantially compromised.
Removing Stage 2: Stage 2 builds upon the groundwork established in Stage 1 to refine the generation of semantically precise and retrieval-efficient query terms, which aims to further augment the quality and utility of retrieved documents. As seen in Table˜2 (1 $\rightarrow$ 3), after removing Stage 2, there is a slight decrease of 5.89% in the accuracy compared to the corresponding model trained with all three stages.
Removing Stage 3: Stage 3 is designed to further enhance the synergy between the model’s retrieval and reasoning capabilities, thereby improving its overall end-to-end performance in solving medical problems. As observed in Table˜2 (1 $\rightarrow$ 2), the performance of models declines by 2.88% when Stage 3 is excluded. However, due to the presence of complete Stage 1 and 2, the gap remains close to that of the complete three-stage training model and relatively low.
C.5 Environment and Hardware Configurations

### v4

@@ B.2. Reward Modeling Details
We provide an in-depth description of the reward computation process associated with progressive RL described in the main content. As shown in the right part of Figure 7, during the verification phase of each training stage, we first parse the model’s natural language output into a structured knowledge graph representation in the form of Equation 1. The stage-specific reward metric is then applied to assess both the natural language response and the structural and logical accuracy of the reasoning trajectory encoded in the knowledge graph.
B.2.1. Stage 1: Reasoner Cultivation
@@
The primary objective of this stage is twofold: (1) to regularize the model’s output format, ensuring structured and consistent responses, and (2) to cultivate its ability to perform logical reasoning when answering medical questions. In addition, by incorporating reward signals based on the number of knowledge retrievals, we encourage the model to make increased use of external medical knowledge sources, which lays the groundwork for enhancing the model’s retrieval capabilities in the subsequent stage. The reward is composed of four aspects: format, answer correctness, reasoning process and retrieval number, as depicted in Figure 7. While format and answer correctness are commonly used reward components, we place particular emphasis on the rationale behind the design of rewards for reasoning process and retrieval number.
@@
  • Alignment with High-Quality Reasoning Trajectories: We aim to guide the model’s reasoning process toward high-quality trajectories observed in frontier models. To this end, our evaluation criteria incorporate semantic and structural characteristics of reference reasoning paths (see Equation 4 to Equation 7), enabling the model to learn from expert-like inference patterns.
  • Capturing Medical Reasoning Specifics: Medical reasoning often requires both logical coherence and comprehensive coverage of relevant entities and their relationships. Unlike mathematical or programming tasks that emphasize rigid logical order, medical reasoning allows for multiple valid pathways, provided they are semantically rich and clinically grounded (Wu et al., 2023; Gao et al., 2025). Accordingly, we introduce: (1) a structural coherence score based on the sequential logic of the reasoning path, as outlined in Equation 7, and (2) a completeness score evaluating the coverage of key entities and relations, as shown in Equation 6.
  • Robust Evaluation via Maximum Similarity Matching: For each training instance, we provide multiple reference reasoning trajectories. However, these references may not be highly similar to one another due to the diversity of valid clinical reasoning. Therefore, in designing Equation 5 to Equation 7, we opt to use the maximum similarity score between the model’s reasoning process and the reference reasoning processes, both in terms of semantics and structure, rather than an average score, to ensure robust and meaningful scoring.
  • Flexibility and Generalization: Leveraging the strengths of reinforcement learning (RL), we avoid rigidly constraining the model’s reasoning trajectory. Instead, we encourage it to approximate high-quality reasoning behaviors while allowing room for variation and creativity. This flexible guidance enhances the model’s generalization capability when faced with novel or out-of-distribution reasoning tasks. To support this goal, we design a multi-dimensional reward scheme that captures various aspects of medical logic reasoning, as depicted in Equation 4.
Table 5. Comparison of average retrieval frequency with and without $\mathcal{R}_{retrieval\_num}$ in Stage 1.
@@
Retrieval Number.  One of the key motivations for incorporating this reward in Stage 1 is to lay the foundation for enhancing the model’s retrieval capabilities in the subsequent stage. Specifically, only when the model develops the habit of querying external knowledge upon encountering unfamiliar concepts during reasoning, can there be meaningful room for optimizing its retrieval behavior. Our experiments also confirm this intuition: without such a reward signal in Stage 1, certain models tend to rely solely on internal knowledge and cease performing external retrievals altogether. As observed in Table 5, using Qwen2.5-7B as an example, the average number of knowledge retrievals during logical reasoning drops significantly when the retrieval count component is omitted from the reward function.
B.2.2. Stage 2: Retriever Awakening
@@
The core reward metric in this phase is the retrieval effectiveness, which evaluates both the quality of the retrieved documents and their influence within the model’s overall reasoning trajectory, as outlined in Equation 9. For document quality assessment, we adopt the Evidence-Based Medicine (EBM) grading framework illustrated in Figure 4 to classify the retrieved documents into different quality levels. To evaluate the contribution of these retrieved documents to the model’s reasoning trajectory, we compute the ratio of retrieved knowledge graph triples relative to the total number of triples in the reasoning trajectory, as defined in Equation 11. This metric reflects the extent to which external knowledge contributes to the logical inference process.
B.2.3. Stage 3: Dual-Process Collaboration
After the previous two stages which have separately enhanced the model’s reasoning and retrieval capabilities, in the final training stage, the model’s end-to-end (E2E) performance on medical problem-solving is directly optimized based on the accuracy of the final answers. As a result, the reward function in this stage consists solely of format and answer correctness.
@@ Appendix C Experiment Details
C.1. Baselines
Here we provide a comprehensive overview of the various models as well as methods that serve as baselines in our comparative analysis.
  • Close-Sourced Models: Close-sourced models are regarded as embodying the current peak performance across various capabilities of LLMs, and serve as the strongest baselines. Here we have selected GPT-4o-mini (Hurst et al., 2024) for comparison.
  • Open-Sourced Medical-Specific Models: These models refer to domain-specific models that were trained specifically on medical data. We have selected MEDITRON-7B, MEDITRON-70B (Chen et al., 2023), UltraMedical3-8B, and UltraMedical3.1-8B (Zhang et al., 2024) to represent the open-sourced medical-specific models as competitors to assess Med-R3’s relative advantage. Specifically, the backbone model of UltraMedical3.1-8B is LLaMA3.1-8B-Instruct, which facilitates a more intuitive comparison with our method.
  • Open-Sourced Medical Reasoning Models: Inspired by the breakthrough of OpenAI o1, these models are capable of complex medical reasoning, applying o1-like methods to the medical field. We have selected HuatuoGPT-o1-8B (Chen et al., 2025a), MedS3-8B (Jiang et al., 2025) and AlphaMed-8B (Liu et al., 2025b) as competitors. HuatuoGPT-o1-8B improved models capabilities in medical reasoning by performing a two-stage SFT + RL training on verifiable medical problems. MedS3-8B equips the model with a self-evolution paradigm and proposes PRM-guided Vote-Sum (P-VS) strategy during inference to enhance long-chain reasoning capabilities in the medical domain. AlphaMed-8B is trained exclusively on rule-based RL. The backbone models of these competitors are all LLaMA-3.1-8B-Instruct, allowing for a more direct comparison with LLaMA-3.1-8B-Instruct + Med-R3.
@@ C.1. Baselines
  • Supervised Fine-Tuning (SFT): The model undergoes SFT using our constructed training dataset, with reasoning trajectories produced by proprietary models (e.g., DeepSeek-V3), enriched with knowledge obtained through retrieval mechanisms. During training, retrieved documents are with masked losses.
  • General Retrieval-Augmented Reasoning RL: We also compare our method with approaches that enhance the model’s retrieval-augmented reasoning capabilities through RL in the general domain. R1-Searcher (Song et al., 2025) is a two-stage outcome-based RL method designed to enhance the model’s capabilities of searching and integrating additional knowledge during the reasoning process. To ensure a more equitable comparison, we replaced the training dataset and the knowledge corpus for retrieval with one that matches our experimental setup. In the first stage, we trained the model for 1 epoch, and in the second stage for 2 epochs, with a total number of training samples matching that of Med-R3. ReSearch (Chen et al., 2025b) considers search operations as integral components of the reasoning chain, and trains models to reason with search via RL without using any supervised data on reasoning steps. Here we also train the model for 3 epochs to align with our experimental configuration. The RL process is implemented by the Group Relative Policy Optimization (GRPO) (Shao et al., 2024) algorithm.
We evaluate these baselines under different inference methods, including Naive (direct generation), CoT (pure reasoning), RAG (pure retrieval), and CoT-RAG (interleaved reasoning and retrieval), as detailed in Table 6.
Table 6. Comprehensive comparison of Med-R3 with baselines, where we have provided the inference strategy for each competitor. Here we have selected the highest-performing one (marked with ) to represent the optimal performance of each method, as summarized in Table 2. * denotes our re-implementation with the same amount of our constructed training data for a fair comparison. represents the PRM guided Vote Sum (P-VS) strategy during inference, which unlocks the full potential of MedS3-8B (Jiang et al., 2025). The best and second best of each model are in bold and underlined.
Model Training Method Inference Method MedQA-USMLE MedQA-MCMLE MedMCQA RareArena-RDC RareArena-RDS MMLU-Med NEJMQA MedXpertQA Avg.
@@
LLaMA3.1-8B -Instruct Naive 31.16 41.45 30.02 41.85 22.16 37.12 50.41 14.90 33.63
@@
C.2. Benchmarks
To evaluate the performance of our proposed Med-R3 in both standard and real-world clinical scenarios of the medical domain, we have selected seven medical datasets, including the MedQA-USMLE, MedQA-MCMLE (Jin et al., 2020), MedMCQA (Pal et al., 2022), MMLU-Med (Hendrycks et al., 2021), RareArena (THUMedInfo, 2025), MedXpertQA (Alonso et al., 2024) and NEJMQA (Katz et al., 2024). Among these benchmarks, MedQA series, MedMCQA and MMLU-Med are standard question-answering tasks, while RareArena, MedXpertQA and NEJMQA are related to real-world clinical scenarios. We employ LLM-as-Judge based on the proprietary model DeepSeek-V3 (Liu et al., 2024) to assess the correctness of the responses, then compute the accuracy scores as our evaluation metric.
We have introduced within-domain tasks in the section of training datasets details, where we use the training subsets for the RL process and test subsets (if no ground-truth labels provided, we use the development subsets) for in-domain evaluation. In the following, we give descriptions of the remaining out-of-domain benchmarks:
  • MMLU-Med (Hendrycks et al., 2021): The MMLU-Med task is composed of six medical domains, anatomy, clinical knowledge, professional medicine, human genetics, college medicine, and college biology, which are extracted from the MMLU benchmark.
  • NEJMQA (Katz et al., 2024): The NEJMQA dataset is constructed from clinical case challenges sourced from The Lancet and the New England Journal of Medicine, which is centered on diagnostic reasoning using patient symptom information.
  • MedXpertQA (Alonso et al., 2024): The MedXpertQA dataset incorporates specialty-specific evaluations and realistic clinical case questions derived from authentic medical practice. Here we utilize the text part of this benchmark.
C.3. More Experimental Observations
The performance of each baseline reported in Table 2 in the main content represents the best result achieved in four different inference strategies. The comprehensive scores across all inference strategies are provided in Table 6. We summarize the additional observations and insights derived from our experiments as follows:
@@ C.3. More Experimental Observations
RL boosts generalizable medical performances. We compare the performance of models after SFT and RL phase. Specifically, R1-Searcher, ReSearch, and Med-R3 all employ RL-based training approaches. As shown in Table 6, on the in-domain benchmarks, the performance gap between SFT and R1-Searcher/ReSearch is relatively small, with SFT underperforming by 0.59% and 5.39%, respectively. However, on out-of-domain tasks, particularly on MMLU-Med, the gap becomes much larger, with SFT lagging behind R1-Searcher and ReSearch by 9.21% and 10.07%, respectively. A primary reason for this discrepancy lies in the nature of SFT as a strongly supervised learning paradigm. Under this setting, the model tends to learn by directly imitating the given reasoning trajectories, effectively memorizing problem-solving patterns rather than truly learning the underlying reasoning process. Consequently, its generalization capability is notably limited when faced with novel or unseen scenarios. It is worth noting that in our SFT setup, the training data includes retrieval-augmented reasoning trajectories that incorporate externally retrieved knowledge, which mitigates the inherent limitations of SFT to some extent. Nevertheless, a clear performance gap still remains.
C.4. Ablation Study and Analysis Details
We report the original performance scores of the models on each benchmark during training stage ablation as well as reward design ablation, as depicted in Table 7 and Table 8, which are the detailed versions of Table 3 and Table 4 in the main content.
Table 7. Original scores for each benchmark of the ablation study on multiple training stages and sequential order. Order: sequential order of Stage 1, 2, and 3. Specifically, “1 &\& 2 &\& 3” refers to a joint training configuration in which reward functions from all three stages are merged and optimized concurrently. The best and second best scores of each model are in bold and underlined. We evaluate these models under the inference strategy of CoT-RAG (interleaved reasoning and retrieval).
@@ C.4. Ablation Study and Analysis Details
1 $\rightarrow$ 2 $\rightarrow$ 3 Qwen2.5-7B 68.64 67.53 68.97 63.02 45.76 75.81 58.54 14.98 57.91
@@
Table 8. Original scores for each benchmark of the ablation study on reward design. w/o: removing the reward component from the original metric. The best and second best scores are marked in bold and underlined. We employ Qwen2.5-7B for experimental analysis, and evaluate the model under the inference strategy of CoT-RAG (interleaved reasoning and retrieval).
@@
C.5. Environment and Hardware Configurations